Circular No. 7952

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 P1 (NEAT) Further observations made of this comet during Aug. 10.4–10.9 UT are

included on MPEC 2002-P46, together with the following very uncertain parabolic orbital elements:

T = 2001 July 7.85 TT  ω = 339.^{o}32
Ω = 312.57 }2000.0
q = 7.0989 AU  i = 43.38

NEAR-SUN COMETS Further to IAUC 7951, D. Hammer reports measurements for comets

detected on C2 SOHO website images by R. Kracht (C/2001 K11 and V6, Meyer group; C/2001 Q8, R8 and R9, Kracht group—see MPEC 2002-E18, 2002-E25, 2002-F03, 2002-O35), M. Oates (C/2001 L10, R7 and X10, Meyer group) and X. Leprette (C/2001 L11, Kreutz group). The reduced astrometry and orbital elements by the undersigned appear on the cited MPEC s.

Comet  2001 UT  α_{2000}  δ_{2000}  MPEC
C/2001 K11 May 16.671  3^{h}37.^{m}3  +20^{o}07′  2002-O34
C/2001 L10 June 1.354  4 39.4  +23 15  2002-O34
C/2001 L11  8.131  5 03.8  +21 03  2002-O35
C/2001 Q8 Aug. 24.868  10 08.3  +12 00  2002-O35
C/2001 R7 Sept. 12.704  11 18.0  + 5 56  2002-O34
C/2001 R8  6.814  10 55.8  + 7 28  2002-O35
C/2001 R9  7.504  10 58.1  + 7 13  2002-O35
C/2001 V6 Nov. 2.329  14 26.0  −13 22  2002-O34
C/2001 X10 Dec. 15.440  17 32.9  −22 04  2002-O34

SUPERNOVA 2002eg IN UGC 11486 A. V. Filippenko, R. Chornock, R. J. Foley, W. D. Li, M. Papenkova

and B. Swift, University of California, Berkeley, report that inspection of CCD spectra (range 310–1000 nm) obtained on Aug. 9 UT with the Shane 3-m reflector at Lick Observatory reveals that SN 2002eg (IAUC 7944) is of type IIb, shortly before maximum brightness. The spectrum is very similar to that of SN 1996cb on 1996 Dec. 27 (Qiu et al. 1997, A.J. 117, 736), ∼ 6 days before B-band maximum; specifically, the Ca II near-infrared triplet is extremely weak or absent, whereas it was present in SN 1996cb at earlier and later epochs.

2002 August 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Brian G. Marsden